# Gaussian Splatting on the Move: Blur and Rolling Shutter Compensation for Natural Camera Motion

### v1

@@ Results
The real data results are given in Table 2 and visualized in Figs. 4 and 5. The pose-optimized version uses the manual calibration data. The corresponding results for the built-in calibration were similar, but slightly worse on the average, see App. 0.D for details.
The motion blur compensation consistently improves the results over the COLMAP baseline, in terms of the PSNR, SSIM and LPIPS metrics, as well as in qualitative evaluation. The effect of the rolling-shutter pose optimization (with non-optimized intrinsics) are mixed. In the Android test sequences, the metrics are clearly and consistently improved, while in the iOS test, they are degraded.
We suspect the main reason for this is the strength of the rolling shutter effect. The tested Andorid devices have long read-out times, $T_{\rm ro}=32.3$ ms and $T_{\rm ro}=28.87$ ms for Pixel 5 and S20, respectively. The readout time of iPhone 15 is approximately 5 ms. The 3DGS pose optimization may be less robust than COLMAP and might more easily converge to wrong local minima. However, in the case of a strong rolling shutter effect, the modeling error from assuming a global shutter is large and COLMAP’s results are easier to improve.
Qualitatively, the sharpness of the reconstructions is subtly, but noticeably increased when motion blur compensation is enabled, as demonstrated in the highlighted details in Figs. 4 and 5. The versions with and without rolling shutter compensation did now show significant differences (cf. App. 0.D). In cases where rolling shutter compensation and pose optimization improved results over the purely motion-blur-compensated case, the improved metrics are mostly explained by more consistent overall shape of the reconstruction, when three-dimensional rolling shutter warping is more accurately modelled.
The automatic selection of key frames for evaluation also played a significant role in our analysis, particularly in highlighting areas poorly represented in the training data. These areas, especially near the edges of the visual field, were more susceptible to artefacts that disproportionately affected the PSNR metric, underlining the challenges in balancing motion blur compensation with the preservation of image quality across the entire scene. These effects are visible near the edges of the error metric figures in Figs. 4 and 5.
Table 3: Real data timing results. Training wall clock time $T$ (minutes) with and without motion blur compensation.
Baseline Motion blur
$N_{\rm train}$ Resolution $T$ $T$
iphone-lego1 21 1920$\times$1440 20 59
iphone-lego2 23 1920$\times$1440 20 91
iphone-lego3 21 1920$\times$1440 45 67
iphone-pots1 25 1920$\times$1440 16 71
iphone-pots2 26 1920$\times$1440 18 50
pixel5-lamp 35 1600$\times$1200 13 65
pixel5-plant 30 1600$\times$1200 12 44
pixel5-table 23 1600$\times$1200 31 37
s20-bike 43 1920$\times$1440 17 51
s20-bikerack 32 1920$\times$1440 18 60
s20-sign 28 1920$\times$1440 20 105
average 21 64
@@
Our results were computed on NVidia A100 GPUs. The training times corresponding to the COLMAP baseline with and without blur are shown in Table 3. The increase in training times (here shown for a single test run) varies significantly by case. On the average, the motion-blur compensated version with $N_{\rm blur}=5$ samples is approximately three times slower than the baseline, which is clearly faster than a 5-fold increase that could be expected in a simpler implementation. However, we note that our CUDA implementation was not yet profiled for new bottlenecks, and could have significant potential for further speed improvement. The memory consumption of our approach is not significantly higher than for the baseline Splatfacto model. For development purposes, we also successfully trained on consumer grade GPUs, such as NVidia RTX 4060 TI with 16 GB of VRAM.
@@ 5 Discussion and Conclusion
We demonstrated how blurring and rolling shutter effects can be efficiently implemented in the Gaussian Splatting (3DGS) framework. We quantitatively and qualitatively demonstrated that the method improves over the 3DGS Splatfacto baseline in both synthetic and real data experiments.
Similar to [25, 18], a relevant additional comparison would be testing against an implementation where training images would be first pre-processed with a deblurring or rolling-shutter unrolling method. However, this paper focused on demonstrating the feasibility of a simpler approach, where the blurring and rolling-shutter effects are directly incorporated into the differentiable 3D implicit representation.
Incorporating learning-based aspects directly into the 3D model generation instead of the 2D input data is a new and promising approach studied in, e.g., [42, 43]. We believe that, in the context of differentiable rendering and 3D reconstruction, this approach is likely to prove superior to learning-based image manipulation as a pre-processing step.
The performance of our approach compared to purely image-based SoTA such as [41, 25] should be analyzed in more detail in follow-up work. Refining VIO-based velocity estimates in 3DGS (or NeRF) during training could also potentially improve reconstruction results. In addition, modelling local linear trajectories with more complex spline-based shapes as in [41, 18] could be an additional follow up work.
@@ 0.A.2 Transforming Gaussians from world to pixel coordinates
where $\tilde{\mu}=K_{i}\hat{\mu}$ and $J_{i}$ is the Jacobian matrix of the pinhole camera projection $\hat{\mu}\mapsto\mu^{\prime}$ with intrisic camera matrix $K_{i}$:
$J_{i}=\left[\begin{matrix}f_{x}&0&0\\ 0&f_{y}&0\end{matrix}\right],\qquad K_{i}=\left[\begin{matrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{matrix}\right].$ (15)
@@ Appendix 0.B Data Sets and Metrics
0.B.1 Modifications to the Deblur-NeRF data set
The original version of the Deblur-NeRF data set [25] did not include the velocity data, which was randomly generated by the original authors with an unspecified seed number. We regenerated the images using the Blender [4] files, which we modified by fixing the random seed, and also included velocity information in our outputs. All of the models mentioned in the [25] were also not available (e.g. ‘trolley’), and those cases are omitted from our version of the data set.
Furthermore, we switched from the custom 10-sample blur implementation to Blender’s built-in motion blur and rolling-shutter effects, which are less prone to sampling artefacts. Finally we adjusted the caustics rendering settings in the Blender scenes to reduce non-deterministic raytracing sampling noise in the data with reasonable sample counts. As a result, our modified data set, published as supplementary material, provides effectively deterministic rendering capabilities and improves the reproducibility of the results compared to the original version.
The modified data set also includes de-focus blur, which was studied in [25] but not in this paper, which is why this variation is also omitted from Table 1.
@@ Appendix 0.C Experiment Details
. The following hyperparameters were tuned in the Splatfacto baseline method to improve obvious issues related to its performance on our datasets: (i) For the synthetic data, we limit the training iteration count to 15k, instead of the default, 30k, to avoid gap-like artefacts that started to appear at higher iteration counts. (ii) For the smarphone data, the iteration count was limited to 20k for the same reason. (iii) The minimum RGB level of all color channels was set to 10 in the training data, to avoid large negative logit color values, which resulted in artefacts in the proximity of dark areas in certain scenes.
The same hyperparameters and modifications were also used for our motion-blur and rolling shutter compensated methods.
@@ Appendix 0.D Additional Results
0.D.1 Synthetic data
The synthetic data results for the remaining scenes are visualized in Fig. 6 and Fig. 7.
@@ 0.D.1 Synthetic data
0.D.2 Smartphone data
The accuracy of the method with manual camera calibration is compared to the built-in calibration values reported by the devices in Table 4, which also includes the results for a method variant where rolling shutter compensation was used without motion blur compensation.
@@ 0.D.2 Smartphone data
A qualitative evaluation of de-blurring with and without rolling shutter compensation is presented in Fig. 8. In both method variants, the details are sharpened in comparison to the baseline.
Finally, the reconstructions for the remaining smartphone datasets presented in Table 2 are shown in Fig. 9.
Table 4: Additional results in the rolling-shutter pose optimization mode: Rolling-shutter only (built-in calibration), and rolling shutter + motion blur compensation with built-in and manual calibration. The rightmost column matches that of Table 2.
RS only (built-in) Built-in calib. Manual calib.
PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$
iphone-lego1 26.40 .859 .193 26.37 .871 .161 26.40 .869 .164
iphone-lego2 26.41 .892 .153 28.13 .907 .133 24.73 .853 .195
iphone-lego3 23.96 .743 .275 24.55 .765 .238 22.93 .720 .282
iphone-pots1 27.66 .864 .204 27.46 .871 .177 27.50 .876 .169
iphone-pots2 25.45 .781 .317 25.83 .806 .255 26.32 .821 .234
pixel5-lamp 28.66 .791 .315 28.11 .793 .293 30.09 .820 .258
pixel5-plant 26.68 .893 .302 26.77 .894 .275 27.32 .908 .242
pixel5-table 29.55 .856 .341 29.91 .858 .312 31.01 .884 .278
s20-bike 27.45 .881 .304 27.92 .885 .298 28.18 .894 .262
s20-bikerack 25.72 .783 .342 25.71 .783 .343 26.68 .797 .297
s20-sign 24.20 .740 .293 24.68 .750 .292 25.89 .780 .259
average 26.56 .826 .276 26.86 .835 .252 27.00 .838 .240
Table 5: Results corresponding to Table 2 with motion-blur-based key frame selection and training-evaluation split disabled. Built-in calibration used in RS pose opt
COLMAP baseline COLMAP + blur RS pose opt + blur
PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$
iphone-lego1 26.68 .890 .200 27.06 .910 .165 26.75 .880 .181
iphone-lego2 27.15 .911 .153 28.01 .925 .135 28.48 .908 .138
iphone-lego3 22.92 .770 .334 23.31 .815 .261 22.62 .780 .281
iphone-pots1 28.18 .912 .211 28.76 .926 .173 27.44 .898 .206
iphone-pots2 28.30 .867 .292 28.63 .881 .242 28.65 .858 .268
pixel5-lamp 27.85 .784 .377 27.81 .784 .347 29.42 .796 .334
pixel5-plant 26.68 .928 .227 26.95 .933 .209 27.85 .929 .239
pixel5-table 28.82 .912 .281 29.34 .922 .260 31.52 .912 .293
s20-bike 26.49 .880 .293 27.64 .889 .271 29.19 .892 .301
s20-bikerack 25.99 .822 .278 26.26 .822 .286 25.33 .787 .367
s20-sign 22.70 .744 .304 23.89 .762 .291 25.26 .766 .303
average 26.52 .856 .268 27.06 .870 .240 27.50 .855 .265
BaselineOurs (MB)Ours (MBRS)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 9: Additional smartphone data reconstructions as in Fig. 4 for the scenes lego3 (iPhone) and sign (S20).

### v2

@@ Results
The real data results are given in Table 3. All features included in the ablation study display a positive impact and our method achieves the best overall performance in therms of the PSNR metric. Performance compared to CVR pre-processing is mixed: while CVR combined with our other compensation features performs best on the data from Pixel 5, it is worse on S20 and clearly worse if enabled for iPhone (with low rolling shutter readout time). Qualitatively we noticed that CVR outputs are often contaminated by various types of artefacts that cause bad performance in 3DGS, and conclude that our method is more robust than CVR for rolling-shutter compensation for static scene reconstructions.
Qualitatively, the sharpness of the reconstructions is subtly, but noticeably increased when motion blur compensation is enabled, as demonstrated in the highlighted details in Figs. 4 and 8. Furthermore, Fig. 5 shows that rolling shutter compensation also noticeably improves the reconstruction quality, especially closer to the edges of the evaluation images.
The automatic selection of key frames for evaluation also played a significant role in our analysis, particularly in highlighting areas poorly represented in the training data. These areas, especially near the edges of the visual field, were more susceptible to artefacts that disproportionately affected the PSNR metric, underlining the challenges in balancing motion blur compensation with the preservation of image quality across the entire scene. These effects are visible near the edges of the error metric figures in Figs. 4 and 8.
Table 4: Real data timing results. Training wall clock time $T$ (minutes) in the baseline method and ours with different features disabled.
[…]
iphone-lego1 21 1920$\times$1440 19 21 81 81
iphone-lego2 23 1920$\times$1440 19 22 78 78
iphone-lego3 21 1920$\times$1440 22 28 114 114
iphone-pots1 25 1920$\times$1440 19 22 76 76
iphone-pots2 26 1920$\times$1440 20 23 83 83
pixel5-lamp 35 1600$\times$1200 16 23 62 67
pixel5-plant 30 1600$\times$1200 15 19 52 58
pixel5-table 23 1600$\times$1200 16 20 55 62
s20-bike 43 1920$\times$1440 21 26 79 90
s20-bikerack 32 1920$\times$1440 20 26 71 79
s20-sign 28 1920$\times$1440 21 36 162 210
average 19 24 83 91
@@
The training times corresponding to selected experiments in Table 3 (which were computed on NVidia A100 GPUs) are shown in Table 4. The increase in training times (here shown for a single test run) varies significantly by case. On the average, our full method with $N_{\rm blur}=5$ samples is approximately five times slower than the baseline. Disabling rolling shutter compensation only has a small, and possibly random effect. However, disabling the motion blur compensation shows that our rolling shutter has a low overhead (26%) compared to the baseline. We also note that our CUDA implementation was not yet profiled for new bottlenecks, and could have significant potential for further speed improvement for motion blur compensation. The memory consumption of our approach is not significantly higher than for the baseline Splatfacto model. For development purposes, we also successfully trained on consumer grade GPUs, such as NVidia RTX 4060 TI with 16 GB of VRAM.
@@ 5 Discussion and Conclusion
We demonstrated how blurring and rolling shutter effects can be efficiently implemented in the Gaussian Splatting (3DGS) framework. We quantitatively and qualitatively demonstrated that the method improves over the 3DGS Splatfacto baseline in both synthetic and real data experiments. Furthermore, we showed that our method outperforms the previous NeRF-based state-of-the-art, BAD-NeRF [42] on synthetic data. We also demonstrated superior performance and robustness compared to baselines where the data was first deblurred or de-rolled using modern deep-learning-based methods, MPR [52] or CVR [6].
Incorporating learning-based aspects directly into the 3D model generation instead of the 2D input data is a new and promising approach studied in, e.g., [43, 44]. We believe that, in the context of differentiable rendering and 3D reconstruction, this approach is likely to prove superior to learning-based image manipulation as a pre-processing step. Modelling local linear trajectories with more complex spline-based shapes as in [42, 18] could be an additional follow up work.
@@ 0.A.2 Transforming Gaussians from world to pixel coordinates
where $\tilde{\mu}=K_{i}\hat{\mu}$ and $J_{i}=J^{\prime}K_{i}$ is the Jacobian matrix of the pinhole camera projection $\hat{\mu}\mapsto\mu^{\prime}$ with intrisic camera matrix $K_{i}$:
[…]
@@ Appendix 0.B Data Sets and Metrics
Modifications to the Deblur-NeRF data set
The original version of the Deblur-NeRF data set [25] (or the BAD-NeRF re-render [42]) did not include the velocity data, which was randomly generated by the original authors with an unspecified seed number. We regenerated the images using the Blender [4] files, which we modified by fixing the random seed, and also included velocity information in our outputs. All of the models mentioned in the [25] were also not available (e.g. ‘trolley’), and those cases are omitted from our version of the data set.
As in [25], the evaluation images in the data set were rendered without blur or rollings shutter effects, which allows us to assess the novel-view synthesis performance of the underlying ‘sharp’ reconstruction independent of the accuracy of our simulation and compensation for these effects.
Furthermore, we switched from the custom 10-sample blur implementation to Blender’s built-in motion blur and rolling-shutter effects, which are less prone to sampling artefacts. In addition, we adjusted the caustics rendering settings in the Blender scenes to reduce non-deterministic raytracing sampling noise in the data with reasonable sample counts. As a result, our modified data set, published as supplementary material, provides effectively deterministic rendering capabilities and improves the reproducibility of the results compared to the original version. Our modified data set also includes de-focus blur, which was studied in [25] but not in this paper, which is why this variation is also omitted from Table 1.
We also note that the different versions of the Deblur-NeRF data set have different gamma correction factors. The original version [25] and our re-render use $\gamma=2.2$ while the BAD-NeRF [42] re-render use $\gamma=1$. In Sec. 4.1, we modify the $\gamma$ parameter of our method accordingly.
Finally, we observed that the Tanabata scene in the BAD-NeRF re-render contains an error that manifests as lower metrics: the sharp evaluation images and blurry training images have been rendered using a slightly different 3D model, apparently unintentionally. This issue is not present in the original version in [25].
@@ Appendix 0.C Experiment Details
The baseline method in our tests is Splatfacto in Nerfstudio version 1.1.0, which includes several improvements from the recent works [50, 51, 45]. For the gsplat code, we used the main branch version accessed on 2024-05-07 and our method has also been implemented based on these versions.
The following non-default parameters were used for both the Splatfacto baseline and our method: (i) limited iteration count to 20k. (ii) enabled antialiased rasterization; (cf. [51]); (iii) enabled scale regularization (cf. [45]). Furthermore, for smartphone data where our method was used with $\gamma=2.2$, the minimum RGB level of all color channels was set to 10 in the training data, to avoid large negative logit color values, which resulted in artefacts in the proximity of dark areas in certain scenes. This modification was only used in our method and not the baseline Splatfacto.
@@ Appendix 0.D Additional Results
Figures
The synthetic data results for the remaining scenes are visualized in Fig. 6 and Fig. 7. The reconstructions for rest of the smartphone data sets presented in Table 3 are shown in Fig. 9.
@@
0.D.0.1 Alternative intrinsics
In addition to the main sequences, we recorded separate manual camera calibration data for smart phones, using the same fixed focus distance as in the other sequences, and used the Kalibr software package (https://github.com/ethz-asl/kalibr) to compute a accurate intrinsic calibration parameters for those devices. We also recorded built-in calibration data reported by the devices.
We then used then trained our method with data where the COLMAP-calibrated intrinsics have been replaced by the manually calibrated intrinsics or the build-in calibration data. The results are presented in Table 5. The built-in calibration data reduced the accuracy in all cases, but the manual calibration improved them for some Android data. We suspect that the reason behind this is that COLMAP also utilizes the camera intrinsic parameters to compensate also for the rolling shutter effect, which results to unoptimal intrinsics solution for a method which can also perform rolling shutter compensation.
We also note that COLMAP-estimated poses do not generally represent the best solution for a given (different) set of camera intrinsics, since slightly inaccurate intrinsics may be compensated by adjusting the poses. Even though this effect can be compensated by our pose optimization, it still appears that for the iPhone data where the rolling shutter effect is negligible, COLMAP-calibrated intrinsics, which are optimized per session, yield better results than manual calibration, which may also have its own inaccuracies.
Blur-based key frame selection
@@ 0.D.0.1 Alternative intrinsics
Table 5: Smartphone results with alternative calibration methods. The COLMAP variant data matches ’Ours’ in Table 3.
COLMAP intrinsics Built-in calibration Manual calibration
@@
iphone-lego1 29.20 .933 .116 27.37 .903 .140 27.37 .902 .143
iphone-lego2 27.95 .926 .117 27.88 .917 .124 27.03 .917 .124
iphone-lego3 24.50 .824 .203 23.58 .759 .250 23.83 .758 .238
iphone-pots1 29.10 .926 .135 28.21 .885 .171 28.25 .883 .177
iphone-pots2 28.00 .882 .180 27.53 .851 .198 27.68 .848 .197
pixel5-lamp 30.46 .833 .196 29.50 .805 .225 30.45 .830 .199
pixel5-plant 27.90 .920 .204 27.14 .902 .239 27.54 .919 .205
pixel5-table 31.86 .919 .196 30.60 .883 .221 32.26 .919 .194
s20-bike 28.93 .911 .242 28.15 .887 .295 29.35 .909 .246
s20-bikerack 29.74 .831 .221 27.62 .814 .269 27.26 .827 .234
s20-sign 26.84 .812 .206 25.31 .773 .249 26.84 .810 .206
average 28.59 .883 .183 27.54 .853 .217 27.99 .866 .197
Table 6: Results corresponding to Table 3 with motion-blur-based key frame selection and training-evaluation split disabled.
Splatfacto Ours
[…]
iphone-lego1 26.26 .892 .202 27.48 .930 .152
iphone-lego2 27.44 .914 .153 27.76 .933 .140
iphone-lego3 24.16 .777 .324 25.63 .853 .198
iphone-pots1 28.39 .917 .202 28.62 .938 .159
iphone-pots2 28.84 .878 .280 29.37 .898 .210
pixel5-lamp 28.36 .865 .336 31.57 .845 .193
pixel5-plant 26.76 .933 .218 28.55 .947 .182
pixel5-table 28.39 .916 .240 31.14 .940 .185
s20-bike 26.81 .904 .288 31.64 .925 .238
s20-bikerack 26.66 .898 .258 30.30 .851 .226
s20-sign 22.78 .782 .290 28.57 .836 .195
average 26.80 .880 .254 29.15 .900 .189
[…]
Figure 9: Additional smartphone data reconstructions as in Fig. 4 for the scenes lego3, pots1 (iPhone) and sign (S20).